OpenStudy (anonymous):

2|x +4|=12 absolute value equations

OpenStudy (anonymous):

1. Isolate the absolute value, so divide 2 on each side. lx+4l=6 2. Now solve for x but there would be 2 equations (6 would be negative and positive) x+4=6 x+4=-6 (x=2 and x=-10)